# Towards Clean-Label Backdoor Attacks in the Physical World

### v1

@@ Abstract
Deep Neural Networks (DNNs) are vulnerable to backdoor poisoning attacks, with most research focusing on digital triggers, special patterns digitally added to test-time inputs to induce targeted misclassification. In contrast, physical triggers, which are natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation. However, current physical attacks require that poisoned inputs have incorrect labels, making them easily detectable upon human inspection. In this paper, we collect a facial dataset of 21,238 images with 7 common accessories as triggers and use it to study the threat of clean-label backdoor attacks in the physical world. Our study reveals two findings. First, the success of physical attacks depends on the poisoning algorithm, physical trigger, and the pair of source-target classes. Second, although clean-label poisoned samples preserve ground-truth labels, their perceptual quality could be seriously degraded due to conspicuous artifacts in the images. Such samples are also vulnerable to statistical filtering methods because they deviate from the distribution of clean samples in the feature space. To address these issues, we propose replacing the standard $\ell_{\infty}$ regularization with a novel pixel regularization and feature regularization that could enhance the imperceptibility of poisoned samples without compromising attack performance. Our study highlights accidental backdoor activations as a key limitation of clean-label physical backdoor attacks. This happens when unintended objects or classes accidentally cause the model to misclassify as the target class.
@@ 1 Introduction
The development of DNNs has led to breakthroughs in various domains, such as computer vision [1], natural language processing [2], speech recognition [3], and recommendation systems [4]. However, training large neural networks requires a huge amount of training data, encouraging practitioners to use third-party datasets, crawl datasets from the Internet, or outsource data collection [5, 6]. These practices introduce a security threat called data poisoning attacks wherein an adversary could poison a portion of training data to manipulate the behaviors of the DNNs [7, 8, 9].
One line of research in data poisoning is backdoor attacks, in which the attackers aim to create an artificial association between a trigger and a target class such that the presence of such trigger in samples from the source class (or multiple source classes) will cause the model to misclassify as the target class. The poisoned model (i.e., the model trained on poisoned samples) behaves normally in ordinary inputs, making backdoor detection challenging. For example, Gu et al. [5] show that a yellow square pattern on a stop traffic sign can change the model’s prediction to “speed limit”.
Most backdoor attacks employ digital triggers, special pixel patterns digitally added at inference time to cause misclassification. Another less explored line of research investigated the use of natural objects within a scene as physical triggers. Physical triggers have two major advantages over digital triggers: they appear more natural when injected into the inputs and do not require digital manipulation during inference, making backdoor attacks on real-time tasks such as facial recognition and object detection possible. However, existing physical backdoor attacks are dirty-label, meaning that the labels of poisoned samples are changed to the label of the target class. The inconsistency between image contents and labels allows data inspectors to easily remove poisoned samples from the target class dataset, especially when such samples contain a visible trigger, such as a special hat, as a common feature. Another issue with current physical backdoor attacks is the unrealistic assumption that attackers can access trigger images from various classes to inject them into the training data, which limits the practicality of physical backdoor attacks in real-world scenarios. Thus, this paper addresses a critical gap: "Is it feasible to execute a successful physical backdoor attack, by introducing a minimal number of clean-label, visually imperceptible poisoned instances into the training set, even when direct access to trigger images is unavailable?"
@@
  • Clean-label: The poisoned inputs preserve the ground-truth labels.
  • Hidden-trigger: The trigger is not present in the poisoned inputs, and the attacker does not need to acquire trigger samples from other classes. Instead, the attacker indirectly plants a backdoor into the model by perturbing a small portion of training data.
  • Real-time: These attacks eliminate the need for digital manipulation at test time, allowing for flexible application in real-time environments, such as facial recognition systems or traffic sign detection for autonomous vehicles.
Currently, there are no public benchmark datasets to study physical backdoor attacks. Therefore, following the design of [10], we collect a large-scale dataset that contains 21,238 images of 10 volunteers and 7 triggers [^1] We followed IRB-approved steps to protect the privacy of our study participants.. Our contributions and findings can be summarized as follows:
Figure 1: General pipeline of clean-label physical backdoor attack. With access to the training dataset and samples with the physical trigger (e.g., sunglasses) from the source class, the attacker uses the attacker model to optimize perturbations that are subsequently added to a small number of target class samples. At inference time, the victim model trained on these perturbed samples will incorrectly classify inputs from the source class with the trigger as belonging to the target class.
@@
    We are the first to conduct a comprehensive empirical study on the performance of clean-label backdoor attacks in the physical world. Our experiments show that the success of physical attacks depends on the poisoning algorithm, physical trigger, and the attacker’s source-target classes. Furthermore, we demonstrate that natural physical backdoors can exist in a dataset if the trigger resembles semantic features of the target class.
@@
    We identify that clean-label attacks can still be susceptible to poison filtering from human inspection and statistical methods. To address this issue, we propose replacing the current $\ell_{\infty}$ regularization with a novel pixel regularization and feature regularization that could enhance the imperceptibility of poisoned inputs in both the pixel and the feature space.
@@
    Our study identifies accidental backdoor activations as a significant limitation of clean-label physical backdoor attacks, occurring when unintended objects or classes cause the model to misclassify as the target class. To mitigate this issue, we introduce repel terms in the adversarial loss that reduce the victim model’s sensitivity to trigger instances from unintended classes.
@@ 2 Related works
In Backdoor Attacks, the attacker uses a predefined trigger to insert a backdoor into a victim model by poisoning a small proportion of training data. Training on the poisoned data causes the model to misclassify instances containing the trigger as the attacker’s chosen target label. Detecting the backdoor attack is challenging because the poisoned model retains high accuracy on clean, ordinary inputs. Generally, backdoor attacks can be divided into dirty-label and clean-label backdoor attacks.
@@
Dirty-label attacks. The attacker could directly enforce a connection between the backdoor trigger and the target class by adding the trigger to the training data and flipping their labels to the target class. Some types of Dirty-label Backdoor Attack include trigger-pattern attacks [5, 11, 12], sample-specific attacks [13, 14], and imperceptible attacks [14, 15, 16, 17, 18]. While dirty-label attacks can achieve impressive performance with a small poisoning ratio, mislabelled poisoned samples are vulnerable to human inspection as their image contents are visibly different from target-class instances.
@@
Clean-label attacks. A more stealthy approach involves directly poisoning target-class instances without label manipulation. The concept of clean-label backdoor attacks was pioneered by Turner et al. [19], who proposed using adversarial perturbations and GAN-based interpolation to obscure the natural, salient features of the target class before embedding the trigger. By effectively concealing the latent features with the perturbations, the model becomes reliant on the introduced trigger for classifying instances of the target class. The following works on Clean-label attacks can be divided into hidden-trigger and trigger-design attacks. In hidden-trigger attacks [20, 21], the trigger is hidden from the training data and only added to test-time inputs of the source class to achieve the targeted misclassification. In trigger-design attacks [22, 23], the attackers aim to optimize trigger pattern(s) that represent the most robust, representative feature of the target class.
@@
Physical-backdoor attacks. Both of these attacks require digitally modifying inputs to insert the trigger. In practical scenarios, however, digital modification is not always possible, especially for real-time applications such as facial recognition and object detection. As an alternative, some works focus on using physical objects as triggers. Notably, Gu et al. [5] demonstrated that a yellow post-it note patched on a stop sign could cause the traffic sign to be misclassified as a speed limit. Similarly, Chen et al. [12] demonstrated an attack that fools a facial recognition system by blending the digital image of sunglasses to the training data and using the physical object (of the same sunglasses) to manipulate the prediction of the model. Wenger et al. [10] conducted an empirical study on the effectiveness of physical objects as backdoor triggers by collecting a dataset with 3,205 images of 9 facial accessories as potential backdoor triggers. These studies on physical backdoor attacks are limited to dirty-label techniques, which is not the focus of our research.
@@
Of particular relevance to our work, Zeng et al. [22] proposed a clean-label backdoor attack that can work in a physical context. However, the lack of details on experimental design, dataset description, and concrete results in their paper hinders any definitive conclusions about the attack’s performance in real-world scenarios. Moreover, their approach differs from our study because it designs a conspicuous trigger pattern instead of exploiting a natural object. By collecting a dataset of 21,238 images with 7 triggers, which is much larger than the dataset of Wenger et al. [10], we are the first to comprehensively study the performance of clean-label backdoor attacks in the physical world.
@@ 3.1 Threat model
Figure 2: Triggers in our dataset (sorted from left to right based on stealthiness)
We further assume that the attacker can select a trigger from a limited pool of physical objects and he has access to trigger source-class samples. This assumption is reasonable given that the source class could be the attacker himself or his collaborators. In some cases, the physical triggers, being natural objects, could already exist in the training dataset of the target class [24] (also referred to as semantic backdoor [25]). In other cases, the attacker who has access to samples from the target class with physical triggers can inject these trigger samples into the training data without raising much suspicion. However, this paper considers a realistic scenario in which the target class dataset does not contain any natural triggers, and the attacker cannot inject trigger instances from the target class.
@@ 3.2 Problem formulation
Notations. Denote $y_{s}$, $y_{t}$ as labels of the source and target classes, respectively. Let $D=\bigcup^{k}_{i=1}D_{i}$ be the training dataset with $k$ classes. Let $\mathcal{D}_{i}$, $\widetilde{\mathcal{D}}_{i}$ be the distribution of instances in class $i$ without and with the physical trigger. Let $F\colon R^{d}\rightarrow R^{k}$ be the classifier model, with parameters $\theta$, that gives prediction scores for each of $k$ classes given an input $x\in R^{d}$. $F$ can be decomposed as $F=f\circ g$ with $f$ as the feature extractor and $g$ be the classifier. The poison pair $s-t$ is the pair of source class and target class with index labels.
Attack goals. The attacker’s goal is to inject a poisoned dataset ${P}$ to the training set ${D}$ such that after training the model on ${D}^{p}:={D}\cup{P}$, it will incorrectly classify samples from the source class wearing the physical trigger as belonging to the target class. To achieve this goal, the attacker needs to optimize the following adversarial objective:
$L_{adv}=\mathbb{E}_{(\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s})}\bigl{[}\mathcal{L}\bigl{(}F_{\theta}(\tilde{x}_{s}),y_{t}\bigr{)}\bigr{]}$ (1)
where $\mathcal{L}$ is the training objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of the source class with the physical trigger. In dirty-label attacks, $L_{adv}$ can be directly optimized by choosing $P$ to be samples drawn from the distribution $\widetilde{\mathcal{D}}_{s}$ and changing the label of each instance in $P$ from $y_{s}$ to $y_{t}$. In clean-label attacks, the attackers indirectly optimize $L_{adv}$ by poisoning samples in $D_{t}$. Thus, we define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data. Existing methods involve crafting a set of perturbations $\delta=\{\delta^{(i)}\}^{K}_{i=1}$ within a $\ell_{\infty}$ norm ball of radius $\epsilon$. These perturbations are then added to a subset of ${D}_{t}$. Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ is the poisoned dataset containing $K$ poisoned target-class examples. In this paper, we introduce four poisoning methods for CLPBA:
  • Naive: Choose $P$ to be a set of samples from $\widetilde{\mathcal{D}}_{t}$ so that the model learns the trigger as features of the target class. This method requires access to trigger instances from the target class.
  • Feature Matching (FM) [20]: This method targets transfer-learning models by finding perturbations that shift poisoned samples closer to the $\widetilde{\mathcal{D}}_{s}$ in the feature space:
    $\mathcal{A}_{FM}=\min_{\|\delta\|_{\infty}\leq\epsilon}\|f(x_{t}+\delta)-f(\tilde{x}_{s})\|_{2}^{2}\quad$ (2)
@@
    $\mathcal{A}_{GM}$ $=\min_{\|\delta\|_{\infty}\leq\epsilon}\left(1-\frac{\nabla_{\theta}L_{poi}\cdot\nabla_{\theta}L_{adv}}{\|\nabla_{\theta}L_{poi}\|\cdot\|\nabla_{\theta}L_{adv}\|}\right)$ (4)
@@
  • Blended Label-Consistent (BLC) [19, 12]: Given an adversarial-trained model, we use the poisoning method in [19] to craft adversarial examples before inserting the digital trigger $p$ (e.g, digital sunglasses) via a perspective transformation $T$ that maps the trigger to appropriate facial area with facial landmarks. The trigger is inserted with an opacity $\tau$, as in [12].
    $x_{t}$ $\leftarrow(1-\tau)\cdot x_{t}+\tau\cdot T(p)$ (7)
    The procedure of BLC is illustrated in Appendix A. Compared to random placement and fixed center placement of the trigger, our adaptive trigger placement method with facial landmarks has a higher attack success rate as it preserves the spatial relationship between the trigger and the face.

### v2

@@ Abstract
Deep Neural Networks (DNNs) are shown to be vulnerable to backdoor poisoning attacks, with most research focusing on digital triggers – special patterns added to test-time inputs to induce targeted misclassification. Physical triggers, natural objects within a physical scene, have emerged as a desirable alternative since they enable real-time backdoor activations without digital manipulation. However, current physical backdoor attacks require poisoned inputs to have incorrect labels, making them easily detectable by human inspection. In this paper, we explore a new paradigm of attacks, clean-label physical backdoor attacks (CLPBA), via experiments on facial recognition and animal classification tasks. Our study reveals that CLPBA could be a serious threat with the right poisoning algorithm and physical trigger. A key finding is that different from digital backdoor attacks which exploit memorization to plant backdoors in deep nets, CLPBA works by embedding the feature of the trigger distribution (i.e., the distribution of trigger samples) to the poisoned images through the perturbations. We also find that representative defenses cannot defend against CLPBA easily since CLPBA fundamentally breaks the core assumptions behind these defenses. Our study highlights accidental backdoor activations as a limitation of CLPBA, happening when unintended objects or classes cause the model to misclassify as the target class. The code and dataset can be found in https://github.com/21thinh/Clean-Label-Physical-Backdoor-Attacks.
@@ 1 Introduction
The development of DNNs has led to breakthroughs in various domains, such as computer vision, natural language processing, speech recognition, and recommendation systems [5, 9, 37, 27]. However, training large neural networks requires a huge amount of training data, encouraging practitioners to use third-party datasets, crawl datasets from the Internet, or outsource data collection [16, 46]. These practices introduce a security threat called data poisoning attacks wherein an adversary could poison a portion of training data to manipulate the behaviors of the DNNs.
One line of research in data poisoning is backdoor attacks, in which the attackers aim to create an artificial association between a trigger and a target class such that the presence of such trigger in samples from the source class, or multiple source classes, will cause the model to misclassify as the target class. The backdoored model (i.e., the model trained on poisoned samples) behaves normally with ordinary inputs while misclassifying trigger instances (i.e., instances injected with the trigger), making backdoor detection challenging. For example, Gu et al. [16] show that a backdoored traffic sign classifier has high accuracy on normal inputs but misclassifies a stop traffic sign as “speed limit” when there is a yellow square pattern on it.
Most backdoor attacks employ digital triggers, special pixel patterns digitally added at inference time to cause misclassification. Another less explored line of research investigated the use of natural objects within a scene as physical triggers. Physical triggers have two major advantages over digital triggers: they appear more natural when injected into the inputs and they do not require digital manipulation during inference. This makes backdoor attacks on real-time tasks such as facial recognition and object detection possible. However, existing physical backdoor attacks are dirty-label, meaning that the labels of trigger samples (i.e., images containing the trigger) are converted to the target-class label. The inconsistency between image contents and labels allows data inspectors to easily remove trigger samples from the dataset, especially when such samples contain a visible trigger (e.g., sunglasses) as a common feature. Thus, this paper addresses the research question: ”Is it feasible to execute a successful physical backdoor attack by injecting a minimal number of clean-label, trigger-free instances with invisible perturbations into the training set?
Figure 1: General pipeline of CLPBA. With access to the training dataset and trigger samples from the source class, the attacker uses the attacker model to optimize perturbations that are subsequently added to a small number of target class samples. At inference time, the victim model trained on these perturbed samples will incorrectly classify the source-class samples with the trigger as the target class.
@@
  • Clean-label: The injected poisoned samples preserve the ground-truth labels.
  • Hidden-trigger: The poisoned samples are perturbed with constrained noise but they do not explicitly contain the trigger. Thus, they are harder to detect since there is no visible difference between poisoned and clean samples.
  • Real-time: These attacks do not require digital manipulation at test time, allowing for applications in real-time environments, such as facial recognition systems or traffic sign detection for autonomous vehicles.
Since there are currently no public benchmark datasets to study physical backdoor attacks, we collected a large-scale facial classification dataset that contains 21,238 images of 10 volunteers and 7 triggers representing common facial accessories, following the design of [55]. We also collected 2000 cat images (500 clean images and 1500 trigger images) containing 3 physical triggers (tennis ball, phone, book) and then combined these images with an existing animal classification dataset [1]. We then conducted extensive experiments on these two datasets to study CLPBA. Our key contributions can be summarized as follows:
@@
    Our study reveals that different from digital backdoor attacks which exploit memorization to plant backdoors in deep nets, CLPBA attacks the model by embedding the features of the source-class trigger distribution (i.e., the distribution of trigger images from the source class that are to be misclassified) to the poisoned images.
@@
    We identify a trade-off between perceptual stealthiness and the effectiveness of CLPBA. To address this issue, we propose a new pixel regularization consisting of a soft version of $\ell_{\infty}$ constraint and Upwind [6] term to smoothen out the perturbations. Empirically, our regularization improves the convergence rate of poisoning algorithms, thereby enhancing the effectiveness and stealthiness of the poisoned samples.
@@
    Our study identifies accidental backdoor activations as a key limitation of CLPBA, occurring when unintended objects or classes cause the model to misclassify as the target class. To address this, we introduce repel terms in the adversarial loss to reduce the victim model’s sensitivity to trigger instances from unintended classes.
  4. 4.
    To facilitate research into physical backdoor attacks, we publicly release the trigger images of the animal classification dataset and the code for reproducibility.
@@ 2 Related Works
In Backdoor Attacks, the attacker uses a predefined trigger to insert a backdoor into a victim model by poisoning a small proportion of training data. Training on the poisoned data causes the model to misclassify instances containing the trigger as the attacker’s chosen target label. Generally, backdoor attacks can be divided into dirty-label and clean-label backdoor attacks.
@@
Dirty-label attacks. The attacker enforces a connection between the backdoor trigger and the target class by adding the trigger to the training data and flipping their labels to the target class. Types of Dirty-label Backdoor Attacks include trigger-pattern attacks [16, 2, 8], sample-specific attacks [39, 30], and imperceptible attacks [30, 11, 12, 38, 34]. While dirty-label attacks achieve impressive performance with a small poisoning ratio, mislabelled poisoned samples are vulnerable to human inspection as their image contents visibly differ from target-class instances.
@@
Clean-label backdoor attacks. A more stealthy approach involves directly poisoning target-class instances without label manipulation. The concept of clean-label backdoor attacks was pioneered by Turner et al. [53], who proposed using adversarial perturbations and GAN-based interpolation to obscure the natural, salient features of the target class before embedding the trigger. By effectively concealing the latent features with the perturbations, the model becomes reliant on the introduced trigger for classifying instances of the target class. The following works on Clean-label attacks can be divided into hidden-trigger and trigger-design attacks. In hidden-trigger attacks [44, 49], the trigger is hidden from the training data and only added to test-time inputs of the source class to achieve the targeted misclassification. In trigger-design attacks [60, 23], the attackers aim to optimize trigger pattern(s) that represent the most robust, representative feature of the target class.
@@
Physical backdoor attacks. Both clean-label and dirty-label backdoor attacks require digitally modifying inputs to insert the trigger. In practical scenarios, however, digital modification is not always possible, especially for real-time applications such as facial recognition and object detection. As an alternative, some works focus on using physical objects as triggers. Notably, Gu et al. [16] demonstrated that a yellow post-it note patched on a stop sign could cause the traffic sign to be misclassified as a speed limit. Similarly, Chen et al. [8] demonstrated an attack that fools a facial recognition system by blending the digital image of sunglasses to the training data and using the physical object (of the same sunglasses) to fool the model at inference. Wenger et al. [55] conducted an empirical study on the effectiveness of physical objects as backdoor triggers by collecting a dataset with 3,205 images of 9 facial accessories as potential backdoor triggers. These studies on physical backdoor attacks are limited to dirty-label techniques, which is not the focus of our research.
@@
Of particular relevance to our work, Zeng et al. [60] proposed a clean-label backdoor attack that can work in a physical context. However, the paper lacks detailed information on experimental design, dataset, and experiments, which hinders definitive conclusions about the attack’s performance in real-world scenarios. Moreover, their approach differs from CLPBA because it designs a conspicuous trigger pattern instead of exploiting a natural object. Thus, we are the first to comprehensively study the performance of clean-label backdoor attacks with physical triggers on two tasks, facial recognition and animal classification.
@@ 3.1 Threat Model
Figure 2: Physical triggers in the collected datasets
We further assume the attacker can select a trigger from a limited pool of physical objects and has access to trigger samples from the source class. This assumption is reasonable, as the source class could be the attacker or collaborators. In some cases, physical triggers, being natural objects, may already exist in the training dataset of the target class [56]. In this paper, we consider a realistic scenario where the target class dataset lacks natural triggers, and the attacker cannot directly inject trigger instances into it.
@@ 3.2 Problem Formulation
For the rest of the paper, we define trigger instances as the clean samples with the injected trigger, and the poisoned instances as the injected samples to the training dataset that backdoors the model. For dirty-label backdoor attacks, poisoned instances are trigger instances with incorrect labels (i.e., mislabeled as the target class). For hidden-trigger backdoor attacks, however, the poisoned instances are images that are perturbed adversarially without the trigger.
Notations. Denote $y_{t}$ as the label of target class. Let $D=\bigcup^{k}_{i=1}D_{i}$ be the training dataset with $k$ classes. Let $\mathcal{D}_{i}$, $\widetilde{\mathcal{D}}_{i}$ be the distribution of instances in class $i$ without and with the physical trigger. Let $F\colon\mathcal{X}\rightarrow R^{k}$ be the classifier model, with parameters $\theta$, that gives prediction scores (i.e., logits) for each of $k$ classes given an input $x\in\mathcal{X}$. $F$ can be decomposed as $F=f\circ g$ with $f$ as the feature extractor and $g$ be the classifier. The poison pair $s-t$ is the pair of source class and target class with index labels.
The attacker aims to inject a poison set ${P}$ to the training set ${D}$ such that the victim model trained on ${D}^{p}:={D}\cup{P}$ will misclassify trigger samples from the source class as the target class. To achieve this goal, the attacker needs to solve the following bilevel optimization problem:
$\underset{\delta\in C}{\text{min}}\;L_{adv}(\delta)=\mathbb{E}_{(\tilde{x}_{s}\sim\widetilde{\mathcal{D}}_{s})}\Bigl{[}\mathcal{L}\bigl{(}F(\tilde{x}_{s};\theta(\delta)),y_{t}\bigr{)}\Bigr{]}$ (1)
$\theta(\delta)\in\operatorname*{arg\,min}_{\theta}\;L_{train}(\theta)=\sum_{x,y\in D^{p}}\mathcal{L}\bigl{(}F(x;\theta),y\bigr{)}$ (2)
where $C$ is the constraint set of the perturbations, $\mathcal{L}$ is the objective function and $\widetilde{\mathcal{D}}_{s}$ is the distribution of trigger instances from the source class. Following previous works [44, 49, 60], the poison set $P$ is a small subset of $D_{t}$. Formally, ${P}=\{x_{t}^{(i)}+\delta^{(i)},y_{t}\}^{K}_{i=1}$ contains $K$ poisoned target-class samples. The constraint set is $\ell_{\infty}$-norm ball: $C=\{\delta\;|\>||\delta||_{\infty}<\epsilon\}$. The value $\epsilon$ controls the magnitude of the perturbations and thus the stealthiness of poisoned samples. We define the poison ratio $\alpha$ as the percentage of poisoned samples in the target-class data. Note that this may differ from previous works on backdoor attacks which define poison ratio over the entire training set. With the above formulation, we introduce three poisoning algorithms for CLPBA based on hidden-trigger backdoor attacks [44, 49] and clean-label data poisoning [21]:
  • Feature Matching (FM) [44]: This method targets transfer-learning models by finding perturbations that shift poisoned samples closer to $\widetilde{\mathcal{D}}_{s}$ in the feature space:
    $\mathcal{A}_{FM}=\min_{\delta\in C}\|f(x_{t}+\delta)-f(\tilde{x}_{s})\|_{2}^{2}\quad$ (3)
@@
  • Meta Poison (MP) [21]: This method is based on meta-learning [13], where it approximates the inner objective $\theta(\delta)$ by unrolling $K$ (default to 2) SGD steps before computing outer objective:
    $\theta_{1}$ $=\theta_{0}-\alpha\nabla_{\theta}L_{train}(\theta_{0}\;|\;\delta_{i})$ (5)
    $\theta_{2}$ $=\theta_{1}-\alpha\nabla_{\theta}L_{train}(\theta_{1}\;|\;\delta_{i})$ (6)
    $\delta_{i+1}$ $=\delta_{i}-\beta\nabla_{\delta}L_{adv}(\delta_{i}\;|\;\theta_{2})\text{ s.t }\delta\in C$ (7)
@@
    $\mathcal{A}_{GM}$ $=\min_{\delta\in C}\left(1-\frac{\nabla_{\theta}L_{poi}\cdot\nabla_{\theta}L_{adv}}{\|\nabla_{\theta}L_{poi}\|\cdot\|\nabla_{\theta}L_{adv}\|}\right)$ (8)
@@
As a baseline, we also compare the poisoning algorithms with the Naive attack, in which the attacker injects trigger target-class samples into the target-class dataset so that the model learns the trigger as a feature of the target class.